Intrigate (1÷logx)-(1÷(logx)^2) upper limit e lower limit 2

Grade:12th pass

1 Answers

Satyam
7 years ago
U could easily break the integration in two parts ;
\int (1).(1/Logx)dx - \int (1/(Logx)^{2})dx
when u would integrate this \int (1).(1/Logx)dx   by using part function(Taking 1 as 2nd function) , it will get cancelled(other part) and u would end up with \int (x/Logx)dx   THEN PUT THE LIMITS AND GET THE ANSWER
 
reach me in future if still have problem
GOOD LUCK!!!